the number of numbers lying between 10 and 1000 can be formed with the digits 2,3,4,0,8,9 is
a)124
b)120
c)125(correct)
d)none of these

Answer by Edwin McCravy(20055)

Case 1: two digit numbers:

Choose the first digit 5 ways (can't choose 0)
Choose the second digit 5 ways (can choose 0)

That's 5*5 = 25 two digit numbers.

Case 2:  three digit numbers

Choose the first digit 5 ways (can't choose 0)
Choose the second digit 5 ways (can choose 0)
Choose the third digit 4 ways.

That's 5*5*4 = 100 three-digit numbers.

Grand total:  25+100 = 125

Edwin